Circular No. 7348

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000C IN NGC 2415 Independent discoveries of a new apparent supernova in NGC 2415

have been made by Steven Foulkes, Ashperton Village, England, on Jan. 8.789 UT (communicated via G. Hurst) and by Marco Migliardi, Cortina, Italy, on Jan. 12.95 (via M. Villi). Foulkes’ discovery was made on an unfiltered CCD image taken with a 0.26-m reflector in the course of the U.K. Nova/Supernova Patrol. Migliardi obtained the following position for SN 2000C from his discovery image: α = 7^{h}36^{m}57^{s}.09, δ = +35^{o}14′38′′.8 (equinox 2000.0), which is 8′′.3 east and 7′′.4 north of the nucleus of NGC 2415. W. Li, University of California at Berkeley, reports that a prediscovery image (limiting mag 19.5–20.0) of SN 2000C appears on a CCD frame taken on Jan. 8.3 with the Katzman Automatic Imaging Telescope (KAIT) at Lick Observatory, from whence he provides the following position end figures: 56^{s}.99, 37′′.2. Li adds that although SN 2000C is very close to the location of SN 1998Y (cf. IAUC 6850), comparison with KAIT images of SN 1998Y shows that the new object is slightly north and east of SN 1998Y

(see http://astron.berkeley.edu/~bait/2000/ngc2415.html); he also

notes that SN 1998Y faded from view on KAIT images in Oct. 1998. Estimated magnitudes for SN 2000C: 1999 Dec. 15.995, [17: (Foulkes); 26, [19.0 (Li); 30, [18.5 (Li); 2000 Jan. 4, 19.5–20.0 (Li; only hint of new object); 8.3, 18.0: (Li); 8.789, 15.6: (Foulkes); 12.95, 15.5 (Migliardi).

COMET C/1999 H3 (LINEAR)

Ephemeris from orbital elements on MPC 35814:

2000 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 7 14^{h}28.^{m}11 +32^{o}31.^{′}3 3.658 3.734 86.^{o}7 15.^{o}2 13.5
17 14 22.59 +33 53.5 3.540 3.766  95.6 15.1 13.5
27 14 14.21 +35 30.6 3.427 3.800 104.8 14.5 13.5
Feb. 6 14 02.47 +37 17.1 3.328 3.836 114.0 13.6 13.4
16 13 46.99 +39 04.6 3.250 3.874 122.7 12.4 13.4
26 13 27.79 +40 41.4 3.200 3.913 130.2 11.1 13.5
Mar. 7 13 05.38 +41 55.0 3.185 3.954 135.4 10.2 13.5
17 12 41.00 +42 34.1 3.210 3.996 137.1  9.8 13.5
27 12 16.41 +42 33.0 3.274 4.039 134.8 10.1 13.6

Visual m_{1} estimates: Jan. 3.80 UT, 13.6 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 6.21, 13.3 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 12.19, 13.1 (C. Segarra, Algimia, Spain, 0.25-m reflector).

2000 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green